Instructions:

Good day! Equations are given below. Please determine which of the following equation is not a homogeneous equation. Thank you!


Choose the letter.


a. (y² - xy) dx + x²dy = 0


b. y' = 2xy e x/y / x2+y² sin(x/y)

Letter b is in fraction form.


c. 3x2ydx + (y+x³) dy = 0


d. 2x2y+3y³ (x³ + 2xy²) y' = 0


1
Expert's answer
2021-06-07T17:34:36-0400

A function f(x,y) is called homogeneous of degree n if f(tx,ty)=tnf(x,y), where t is independent of x and y and t belongs to real numbers.

Option (c) is not homogeneous.

"3x^2ydx+(y+x^2)dy=0\\
y'=\frac{-3x^2y}{x^3+y}\\
put 
x=tX and y=tY, we \space get\\
Y'=\frac{t^3(-3X^2Y)}{t(t^2X^3+Y)}\neq t^nf(x,y)\\
\text{And other options are homogeneous.}"

bsns